# Dynamic Subgraph Matching via Cost-Model-based Vertex Dominance Embeddings (Technical Report)

### v2

@@ Abstract.
In many real-world applications such as social network analysis, knowledge graph discovery, biological network analytics, and so on, graph data management has become increasingly important and has drawn much attention from the database community. While many graphs (e.g., Twitter, Wikipedia, etc.) are usually involving over time, it is of great importance to study the dynamic subgraph matching (DSM) problem, a fundamental yet challenging graph operator, which continuously monitors subgraph matching results over dynamic graphs with a stream of edge updates. To efficiently tackle the DSM problem, we carefully design a novel vertex dominance embedding approach, which effectively encodes vertex labels that can be incrementally maintained upon graph updates. Inspire by low pruning power for high-degree vertices, we propose a new degree grouping technique over basic subgraph patterns in different degree groups (i.e., groups of star substructures), and devise degree-aware star substructure synopses (DAS3) to effectively facilitate our designed vertex dominance and range pruning strategies. We develop efficient algorithms to incrementally maintain dynamic graphs and answer DSM queries. Through extensive experiments, we confirm the efficiency of our proposed approaches over both real and synthetic graphs.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/JamesWhiteSnow/DSM.
@@ 1. Introduction
While many previous works (He and Singh, 2008; Shang et al., 2008; Bonnici et al., 2013; Bi et al., 2016; Jüttner and Madarasi, 2018; Han et al., 2019; Bhattarai et al., 2019; Sun and Luo, 2020) usually considered the subgraph matching over static graphs, real-world graph data are often dynamic and change over time. For example, in social networks, friend relationships among users may be subject to changes (e.g., adding or breaking up with friends); similarly, in collaboration networks, people may start to collaborate on some project and then suddenly stop the collaboration for years. In these scenarios, it is important, yet challenging, to conduct the subgraph matching over such a large-scale, dynamic data graph $G_{D}$, upon updates (e.g., edge insertions/deletions). In other words, we need to continuously monitor subgraphs (e.g., user communities or collaborative teams) in $G_{D}$ that follow query graph patterns $q$.
(a) query graph $q$
(b) dynamic collaboration network $G_{D}$
Figure 1. An example of the subgraph matching in dynamic collaboration networks $G_{D}$.
@@
Figure 1(b) provides an example of a dynamic collaboration network $G_{D}$, where edges are inserted or deleted over time. For example, a new edge $(v_{4},v_{7})$ is inserted at timestamp $t=1$, implying that two experts $v_{4}$ and $v_{7}$ start to collaborate with each other in a team. Similarly, at timestamp $t=2$, an existing edge $(v_{6},v_{7})$ is removed from the graph $G_{D}$, which may indicate that experts $v_{6}$ and $v_{7}$ have not collaborated with each other for a certain period of time (e.g., two years).
@@
In this example, the job search advisor can register a dynamic subgraph matching query over $G_{D}$ to continuously identify subgraphs of $G_{D}$ that match with the query graph pattern $q$. At timestamp $t=1$, subgraphs $g_{1}$ and $g_{2}$ are the subgraph matching answers, whereas at timestamp $t=2$, subgraph $g_{1}$ is the only answer (since the deletion of edge $(v_{6},v_{7})$ invalidates the subgraph $g_{2}$).  $\blacksquare$
@@
Dynamic subgraph matching has many other real applications. For example, in the application of anomaly detection for online shopping (Qiu et al., 2018), the interactions (e.g., purchases, views, likes, etc.) among users/products over time form a dynamic social/transaction graph, and it is important to detect some abnormal events such as fraudulent activities (i.e., query graph patterns) over such a dynamically changing graph.
@@
Inspired by the examples above, we formulate a dynamic subgraph matching (DSM) query over a dynamic graph, which continuously monitors subgraph matching results, upon updates to the data graph.
Prior Works: Due to NP-completeness of the subgraph isomorphism (Garey and Johnson, 1983; Cordella et al., 2004; Grohe and Schweitzer, 2020), dynamic subgraph matching is not tractable. To tackle this problem, several exact algorithms have been proposed (Fan et al., 2013b; Choudhury et al., 2015; Kankanamge et al., 2017; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021), which consist of three categories: recomputation-based (Fan et al., 2013b), direct-incremental (Kankanamge et al., 2017), and index-based incremental algorithms (Choudhury et al., 2015; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021). The recomputation-based algorithms re-compute all the subgraph matching answers at each timestamp from scratch, the direct-incremental algorithms incrementally calculate new matching results from dynamic graphs upon updates, and index-based incremental ones incrementally maintain matching results via auxiliary indexes built over query answers.
@@
Our Contributions: Our work falls into the category of the direct-incremental algorithm. Different from existing works that directly use structural information (e.g., vertex label and neighbors’ label set) to filter out vertex candidates, we design a novel and effective vertex dominance embedding technique for candidate vertex/subgraph retrieval. Specifically, our proposed vertex dominance embeddings can be incrementally maintained in dynamic graphs, and transform our DSM problem into a dominating region search problem in the embedding space with no false dismissals. To enhance the pruning power of our embeddings for high-degree vertices, we propose a new degree grouping approach for vertex embeddings over basic subgraph patterns in different degree groups (i.e., groups of star substructures). We also devise a cost model to further guide the vertex embeddings to effectively achieve high pruning power. Finally, we develop algorithms to incrementally maintain dynamic graphs and answer DSM queries efficiently.
Specifically, in this paper, we make the following contributions:
@@
    We formally define the dynamic subgraph matching (DSM) problem in Section 2, and propose a general framework for DSM query answering in Section 3.
@@
    We carefully design incremental vertex dominance embeddings to facilitate dynamic subgraph matching in Section 4.
@@
    We devise a degree grouping technique to enhance the pruning power of vertex embeddings, and construct degree-aware star substructure synopses (DAS3) to support the dynamic subgraph matching in Section 5.
@@
    We design effective synopsis pruning strategies, and develop efficient DSM processing algorithms in Section 6.
@@
    We present a novel cost-model-guided vertex embeddings to further improve the pruning power in Section 7.
@@
    We demonstrate through extensive experiments the efficiency and effectiveness of our DSM processing algorithm over real/synthetic graphs in Section 8.
@@
2. Problem Definition
Table 1 depicts the commonly used symbols in this paper.
2.1. Static Graph Model
We give the model of a static undirected, vertex-labeled graph $G$.
@@ 2.1. Static Graph Model
[…]
The static graph model (as given in Definition 2.1) has been widely used in various real-life applications to reflect relationships between different entities, for example, social networks (Al-Baghdadi et al., 2020), Semantic Web (Hassanzadeh et al., 2012), transportation networks (Rai and Lian, 2023), biological networks (Kan et al., 2023), citation networks (Yang and Han, 2023), and so on.
Definition 0.
In Definition 2.2, the graph isomorphism problem checks whether graphs $G_{1}$ and $G_{2}$ exactly match with each other.
Definition 0.
Note that, the subgraph matching problem is NP-complete (Garey and Johnson, 1983).
2.2. Dynamic Graph Model
Real-world graphs are often continuously evolving and updated. Examples of such dynamic graphs include social networks with updates of friend relationships among users, bibliographical networks with new collaboration relationships, and so on. In this subsection, we provide the data model for dynamic graphs with update operations below.
Definition 0.
[…]
Snapshot Graph, $G_{t}$: From Definition 2.4, after applying to initial graph $G$ all the graph update operations up to the current timestamp $t$ (i.e., $\Delta G_{1}$, $\Delta G_{2}$, $\dots$, and $\Delta G_{t}$), we can obtain a snapshot of the dynamic graph $G_{D}$, denoted as $G_{t}$.
@@ 2.2. Dynamic Graph Model
We will later discuss how to deal with the three cases of edge insertions above for dynamic graph maintenance and incremental query answering.
2.3. Dynamic Subgraph Matching Queries
In this subsection, we define a dynamic subgraph matching (DSM) query in a large dynamic graph $G_{D}$, which continuously monitors subgraph matching answers for a set, $Q$, of query graph patterns.
@@ 2.3. Dynamic Subgraph Matching Queries
(Dynamic Subgraph Matching, DSM) Given a dynamic graph $G_{D}$, a set, $Q$, of query graph patterns $q$, and a timestamp $t$, a dynamic subgraph match (DSM) query maintains a subgraph matching answer set $A(q,t)$ for each query graph pattern $q\in Q$, such that any subgraph $g\in A(q,t)$ in $G_{D}$ is isomorphic to the query graph $q$ (denoted as $g\equiv q$).
Alternatively, for each $q\in Q$, the DSM problem incrementally computes an update, $\Delta A(q,t)$, to the answer set $A(q,t-1)$, upon the change $\Delta G_{t}$ to $G_{t-1}$. $\blacksquare$
In Definition 2.5, the DSM query continuously maintains a subgraph matching answer set $A(q,t)$ (w.r.t. each query graph pattern $q\in Q$) over dynamic graph $G_{D}$ (upon incremental updates) for a long period of time. The DSM problem is quite useful in real applications such as cyber-attack event detection and credit card fraud monitoring.
@@ 2.4. Challenges
The subgraph matching problem in a data graph has been proven to be NP-complete (Garey and Johnson, 1983). Thus, it is even more challenging to design effective techniques for answering subgraph matching queries in the scenario of dynamic graphs. Due to frequent updates to the data graph, it is non-trivial how to effectively maintain a large-scale dynamic graph that can support effectively obtaining initial subgraph matching results, and continuously and efficiently monitor DSM answer sets, upon fast graph updates.
To tackle the DSM problem, in this paper, we will propose a novel and effective vertex dominance embedding technique over dynamic graphs, which transforms the subgraph matching to a dominance search problem in a graph-node-embedding space. Our proposed embedding technique can guarantee the subgraph matching with no false dismissals. Furthermore, we also design effective pruning, indexing, and incremental maintenance mechanisms to facilitate efficient and scalable dynamic subgraph matching.
@@
// Graph Maintenance Stage
@@
3generate a vertex dominance embedding vector $o(v_{i})$ for each vertex $v_{i}\in G_{0}$
@@
5build $m$ synopses, $Syn_{j}$, over vertex embeddings with degree grouping
@@
15for each query graph $q\in Q$ do
@@
33            apply changes $\Delta A(q,t)$ to $A(q,t-1)$ and obtain the latest subgraph answers in $A(q,t)$
@@
35            return $A(q,t)$
36
37
Algorithm 1 The Dynamic Subgraph Matching Framework
3. The Dynamic Subgraph Matching Framework
In this section, we illustrate a general framework for dynamic subgraph matching in Algorithm 1, which consists of two stages, graph maintenance and DSM query answering stages.
Specifically, in the graph maintenance stage, we first pre-process initial graph $G_{0}$, by constructing a set of $m$ graph synopses, $Syn_{j}$, over vertex dominance embeddings $o(v_{i})$ with $m$ degree groups, respectively (lines 1-2). Then, for each graph update operation $\Delta G_{t}\in\Delta\mathcal{G}$, we incrementally maintain the dynamic graph $G_{t}$ and synopses $Syn_{j}$ (lines 3-5).
In the DSM query answering stage, for each newly registered query graph $q$, we obtain the initial query answer set over the current snapshot $G_{t}$ of the dynamic graph $G_{D}$ (lines 6-10). Then, we monitor the updates in query answer sets over continuously changing dynamic graph $G_{t}$ (lines 11-16).
In particular, for initial answer set generation, we compute query vertex embeddings $o(q_{i})$ from vertices $q_{i}\in V(q)$ in each query graph $q$ (lines 6-7). Then, for each query vertex $q_{i}$, we find and store candidate vertices in a set $q_{i}.cand\_set$ by accessing $Syn_{j}$ (line 8). After that, we enumerate candidate subgraphs $g$ by assembling candidate vertices in $q_{i}.cand\_set$ (line 9), and refine/return matching subgraphs $g$ in an initial subgraph matching answer set $A(q,t)$ (line 10).
Next, to answer the DSM query, for each graph update operation $\Delta G_{t}\in\Delta\mathcal{G}$ and each query graph $q\in Q$, we check the updated vertices in $\Delta G_{t}$, update candidate vertex sets $q_{i}.cand\_set$, and monitor the changes $\Delta A(q,t)$ in $A(q,t)$ (lines 11-14). Finally, we obtain $A(q,t)$ by applying changes $\Delta A(q,t)$ to $A(q,t-1)$, and return $A(q,t)$ as the DSM query answer set (lines 15-16).
4. Incremental Vertex Dominance Embeddings
In this section, we will present our vertex dominance embeddings, which can be incrementally maintained, preserve dominance relationships between basic subgraph patterns (i.e., star subgraphs), and support efficient subgraph matching over dynamic graphs.

### v3

# Continuous Subgraph Matching via Cost-Model-based Dynamic Vertex Dominance Embeddings (Technical Report)

@@ Abstract.
In many real-world applications such as social network analysis, knowledge graph discovery, biological network analytics, and so on, graph data management has become increasingly important and has drawn much attention from the database community. While many graphs (e.g., Twitter, Wikipedia, etc.) are usually evolving over time, it is of great importance to study the continuous subgraph matching (CSM) problem, a fundamental, yet challenging, graph operator, which continuously monitors subgraph matching results over dynamic graphs with a stream of edge updates. To efficiently tackle the CSM problem, we carefully design a general CSM processing framework, based on novel DynamIc Vertex DomINance Embedding (DIVINE), which maps vertex neighborhoods into an embedding space to enable efficient subgraph matching and incremental maintenance under dynamic updates. Inspired by low pruning power for high-degree vertices, we propose a new degree grouping technique to decompose high-degree star patterns into groups of lower-degree star substructures, and devise degree-aware star substructure synopses (DAS3) over embeddings of star substructure groups. We develop efficient algorithms to incrementally maintain dynamic graphs and answer CSM queries by traversing DAS3 synopses and applying our designed vertex dominance and range pruning strategies. Through extensive experiments, we confirm the efficiency of our proposed DIVINE approach over both real and synthetic graphs.
Vertex Dominance Embedding, Continuous Subgraph Matching
copyright: acmlicensedjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 6 (SIGMOD)article: 309publicationmonth: 12doi: 10.1145/3769774ccs: Data Models and Languages Graphs, social networks, web data, and semantic web
@@ 1. Introduction
While many previous works (He and Singh, 2008; Shang et al., 2008; Bonnici et al., 2013; Bi et al., 2016; Jüttner and Madarasi, 2018; Han et al., 2019; Bhattarai et al., 2019; Sun and Luo, 2020) usually considered the subgraph matching over static graphs, real-world graph data are often dynamic and changing over time. For example, in social networks, friend relationships among users may be subject to changes (e.g., adding or breaking up with friends); similarly, in collaboration networks, people may start to collaborate on some project and then suddenly stop the collaboration for years. In these scenarios, it is important, yet challenging, to conduct the subgraph matching over such a large-scale, dynamic data graph $G_{D}$, upon updates (e.g., edge insertions/deletions). In other words, we need to continuously monitor subgraphs (e.g., user communities or collaborative teams) in $G_{D}$ that follow query graph patterns $q$ in real-world applications (such as online advertising to communities on social networks or finding collaboration teams from bibliographic networks).
(a) An example of the subgraph matching in dynamic collaboration networks $G_{D}$.
@@
Figure LABEL:subfig:initial_graph provides a dynamic collaboration network $G_{D}$, where collaboration edges among experts are inserted or deleted over time. For example, a new edge $(v_{4},v_{7})$ is inserted at timestamp $t_{1}=100$, implying that two experts $v_{4}$ and $v_{7}$ start to collaborate with each other in a team. Similarly, at timestamp $t_{2}=300$, an existing edge $(v_{6},v_{7})$ is removed from the graph $G_{D}$, which may indicate that experts $v_{6}$ and $v_{7}$ have not collaborated with each other for a certain period of time (e.g., two years).
@@
In this example, the job search advisor can register a continuous subgraph matching query over $G_{D}$ to continuously identify subgraphs of $G_{D}$ that match with the query graph pattern $q$. At timestamp $t_{1}$, subgraphs $g_{1}$ and $g_{2}$ are the subgraph matching answers, whereas at timestamp $t_{2}$, subgraph $g_{1}$ is the only answer (since the deletion of edge $(v_{6},v_{7})$ invalidates the subgraph $g_{2}$).  $\blacksquare$
@@
Continuous subgraph matching over dynamic graph has many other real applications. For example, in the application of anomaly detection for online shopping (Qiu et al., 2018), the interactions (e.g., purchases, views, likes, etc.) among users/products over time form a dynamic social/transaction graph, and it is important to detect some abnormal events or fraudulent activities (i.e., query graph patterns) over such a dynamically changing graph.
@@
Inspired by the examples above, we formulate a continuous subgraph matching (CSM) query over a dynamic graph, which continuously monitors subgraph matching results, upon updates to the data graph.
Prior Works: Existing exact continuous matching methods obtain incremental matching results in the following three ways: i) recomputing matching results at each timestamp (Fan et al., 2013b); ii) direct updating matching results over the data graph (Kankanamge et al., 2017), and; iii) using auxiliary index over matching results to find updates (Choudhury et al., 2015; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021).
@@
Our Contributions: Different from existing works that directly use structural information (e.g., vertex label and neighbors’ label set) to filter out vertex candidates, In this paper, we design a novel and effective DynamIc Vertex DomINance Embedding (DIVINE) approach for candidate vertex/subgraph retrieval. Specifically, we propose a dynamic vertex dominance embedding technique, which transforms our CSM problem to a dominating region search problem in the embedding space with no false negatives, and allows the generated vertex embeddings to be incrementally maintained in dynamic graphs. To enhance the pruning power of our embeddings for high-degree vertices, we propose a new degree grouping approach for vertex embeddings over basic subgraph patterns in different degree groups (i.e., groups of star substructures). We also devise a cost model to further guide the vertex embedding generation to effectively achieve high pruning power. Finally, we develop algorithms to incrementally maintain dynamic graphs and answer CSM queries efficiently.
Specifically, we make the following main contributions:
@@
    We formally define the continuous subgraph matching (CSM) problem in Section 2, and propose a general framework for CSM query answering in Section 3.
@@
    We carefully design incremental vertex dominance embeddings to facilitate CSM processing in Section 4.
@@
    We devise a degree grouping technique to enhance the pruning power of vertex embeddings, and construct degree-aware star substructure synopses (DAS3) to support the continuous subgraph matching in Section 5.
@@
    We design effective synopsis pruning strategies, and develop an efficient algorithm, named DynamIc Vertex DomINance Embedding (DIVINE), for CSM processing in Section 6.
@@
    We present novel cost-model-guided vertex embeddings to further improve the CSM pruning power in Section 7.
@@
    We demonstrate through extensive experiments the efficiency and effectiveness of our DIVINE approach over real/synthetic graphs in Section 8.
@@
2. Problem Definition
Table 1 depicts the commonly used symbols and their descriptions in this paper.
2.1. Static Graph Model
We give the model of a static undirected, vertex-labeled graph $G$.
Definition 0.
(Static Graph, $G$) A static graph, $G$, is denoted as a triple $(V(G),E(G),L(G))$, where $V(G)$ is a set of vertices $v_{i}$, $E(G)$ is a set of edges $e_{i,j}=(v_{i},v_{j})$ between vertices $v_{i}$ and $v_{j}$, and $L(G)$ is a label function from each vertex $v_{i}\in V(G)$ to a label $l(v_{i})$.  $\blacksquare$
The static graph model (as given in Definition 2.1) has been widely used in various real-life applications to reflect relationships between different entities, such as social networks (Al-Baghdadi et al., 2020), Semantic Web (Hassanzadeh et al., 2012), transportation networks (Rai and Lian, 2023), biological networks (Kan et al., 2023), citation networks (Yang and Han, 2023), and so on.
Definition 0.
In Definition 2.2, the graph isomorphism problem checks whether the graphs $G_{1}$ and $G_{2}$ exactly match each other.
Definition 0.
Note that, the subgraph matching problem has been proven to be NP-complete (Garey and Johnson, 1983).
@@ 2.1. Static Graph Model
2.2. Dynamic Graph Model
Real-world graphs are often continuously evolving and updated. Examples of such dynamic graphs include social networks with new or broken friend relationships among users, bibliographical networks with new/discontinued collaboration relationships, and so on. In this subsection, we provide the data model for dynamic graphs with update operations (i.e., edge insertion/deletion) below.
@@ 2.2. Dynamic Graph Model
(Dynamic Graph, $G_{D}$) A dynamic graph, $G_{D}$, consists of an initial graph $G_{0}$ (as given by Definition 2.1) and a sequence of graph update operations $\Delta\mathcal{G}=\{\Delta G_{1},\Delta G_{2},\dots,\Delta G_{t},\dots\}$ on $G$. Here, $\Delta G_{t}$ is a graph update operator, in the form of $(e+,t)$ or (e,t)(e-,t), which indicates either an insertion (++) or a deletion (-) of an edge $e$ at timestamp $t$, respectively.   $\blacksquare$
Note that, our dynamic graph model follows the widely used edge-centric update paradigm (Choudhury et al., 2015; Kankanamge et al., 2017; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021; Sun et al., 2022a), which has many real-world dynamic graph applications, such as adding/deleting social relationships or collaborations over time. In this model, the insertion of an edge may implicitly introduce new vertices (i.e., ending vertices of the edge), and the deletion of an edge will not remove (isolated) vertices. Thus, the graph size may monotonically grow, as our dynamic graph model evolves. We would like to study topics of supporting isolated vertex insertions/deletions in dynamic graphs as our future work.
Snapshot Graph, $G_{t}$: From Definition 2.4, after applying to initial graph $G_{0}$ all the graph update operations up to the current timestamp $t$ (i.e., $\Delta G_{1}$, $\Delta G_{2}$, $\dots$, and $\Delta G_{t}$), we can obtain a snapshot of the dynamic graph $G_{D}$, denoted as $G_{t}$.
@@
We will later discuss how to deal with these three cases of edge insertions above for dynamic graph maintenance and incremental query answering.
2.3. Continuous Subgraph Matching Queries
In this subsection, we define a continuous subgraph matching (CSM) query in a large dynamic graph $G_{D}$, which continuously monitors subgraph matching answers for a set, $Q$, of query graph patterns.
@@ 2.3. Continuous Subgraph Matching Queries
(Continuous Subgraph Matching, CSM) Given a dynamic graph $G_{D}$, a set, $Q$, of registered query graph patterns $q$, and a current timestamp $t$, a continuous subgraph matching (CSM) query maintains a subgraph matching answer set $A(q,t)$ for each query graph pattern $q\in Q$, such that any subgraph $g\in A(q,t)$ in $G_{D}$ is isomorphic to the query graph $q$ (denoted as $g\equiv q$).
Alternatively, for each $q\in Q$, the CSM problem incrementally computes an update, $\Delta A(q,t)$, to the answer set A(q,t1)A(q,t-1), upon the change $\Delta G_{t}$ to $G_{t-1}$. $\blacksquare$
The CSM problem in Definition 2.5 is useful in real applications such as cyber-attack event detection in computer networks and credit card fraud monitoring over transaction networks. In particular, CSM incrementally maintains the subgraph matching answer set, $A(q,t)$, which contains subgraphs $g\subseteq G_{t}$ isomorphic to query graph $q$ at current timestamp $t$, where $t$ is the current logical time or update step of dynamic graph $G_{D}$. Note that, the answer set $A(q,t)$ may change for different timestamps $t$, upon edge insertions/deletions. For example, $A(q,2)$ and $A(q,5)$ can be different, due to graph updates $\Delta G_{3}$, $\Delta G_{4}$, and $\Delta G_{5}$.
@@ 2.4. Challenges
The subgraph matching problem in a data graph has been proven to be NP-complete (Garey and Johnson, 1983). Thus, it is even more challenging to design efficient and effective techniques for answering subgraph matching queries in the scenario of dynamic graphs. Due to frequent updates to the data graph, it is non-trivial how to effectively maintain a large-scale dynamic graph that can support efficiently obtaining initial subgraph matching results, and continuously monitor CSM answer sets, upon fast graph updates.
@@
// Graph Maintenance Phase
@@
3generate a vertex embedding $o(v_{i})$ for each vertex $v_{i}\in G_{0}$
@@
5build $m$ synopses, $Syn_{j}$, over vertex dominance embeddings with degree grouping
@@
15for each new query graph $q\in Q$ do
@@
33     apply changes $\Delta A(q,t)$ to A(q,t1)A(q,t-1) and obtain the latest subgraph answers in $A(q,t)$
@@
35
Algorithm 1 The DIVINE-based Framework for Continuous Subgraph Matching
3. Continuous Subgraph Matching Framework
Algorithm 1 illustrates a general framework for continuous subgraph matching based on DynamIc Vertex DomINance Embedding (DIVINE), which consists of two phases, graph maintenance and CSM query answering phases. That is, we first pre-process the initial graph $G_{0}$ by constructing $m$ synopses $Syn_{j}$ over our proposed dynamic vertex dominance embeddings, and incrementally maintain the update (lines 1-5). Here, the subscript $j$ in $Syn_{j}$ refers to the synopsis structure corresponding to the $j$-th degree group of vertices. Then, we answer the CSM query over synopses via initial answer generation (lines 6-10) and dynamic answer monitoring (lines 11-15).
Specifically, in the graph maintenance phase, we first pre-process the initial graph $G_{0}$, by constructing $m$ graph synopses, $Syn_{j}$, over vertex dominance embeddings $o(v_{i})$ with $m$ degree groups (discussed later in Section 5), respectively (lines 1-2). Here, the subscript $j$ in $Syn_{j}$ refers to the synopsis structure corresponding to the $j$-th degree group of vertices. Then, for each graph update operation $\Delta G_{t}\in\Delta\mathcal{G}$, we incrementally maintain the dynamic graph $G_{t}$, vertex embeddings, and synopses $Syn_{j}$ (lines 3-5).
In the CSM query answering phase, for each newly registered query graph $q$, we obtain the initial query answer set, $A(q,t)$, over the current snapshot $G_{t}$ of dynamic graph $G_{D}$ (lines 6-10). Then, we monitor the updates in query answer sets (w.r.t. $Q$) over continuously changing dynamic graph $G_{t}$ (lines 11-15).
In particular, for initial answer set generation, we first compute query vertex embeddings $o(q_{i})$ from vertices $q_{i}\in V(q)$ in each query graph $q$ (lines 6-7). Then, for each query vertex $q_{i}$, we find and store candidate vertices in a set $q_{i}.cand\_set$ by accessing $Syn_{j}$ (line 8). After that, we enumerate candidate subgraphs $g$ by assembling candidate vertices in $q_{i}.cand\_set$ (line 9), and refine/return matching subgraphs $g$ in an initial subgraph matching answer set $A(q,t)$ (line 10).
Next, to monitor CSM query answers, for each graph update operation $\Delta G_{t}\in\Delta\mathcal{G}$ and each query graph $q\in Q$, we check the updated vertices in $\Delta G_{t}$, update candidate vertex sets $q_{i}.cand\_set$, and compute the changes $\Delta A(q,t)$ in A(q,t1)A(q,t-1) (lines 11-14). Finally, we obtain the CSM query answer set $A(q,t)$ by applying changes $\Delta A(q,t)$ to A(q,t1)A(q,t-1) (line 15).
4. Dynamic Vertex Dominance Embeddings
In this section, we will present our dynamic vertex dominance embeddings, which can be incrementally maintained, preserve dominance relationships between basic subgraph patterns (i.e., unit star subgraphs), and support efficient CSM over dynamic graphs.